### v1

# Intriguing low-temperature phase in the antiferromagnetic kagome metal FeGe

Abstract
The properties of kagome metals are governed by the interdependence of band topology and electronic correlations resulting in remarkably rich phase diagrams. Here, we study the temperature evolution of the bulk electronic structure of the antiferromagnetic kagome metal FeGe using infrared spectroscopy. We uncover drastic changes in the low-energy interband absorption at the 100 K structural phase transition that has been linked to a charge-density-wave (CDW) instability. We explain this effect by the minuscule Fe displacement in the kagome plane, which results in parallel bands in the vicinity of the Fermi level. In contrast to conventional CDW materials, however, the spectral weight shifts to low energies, ruling out the opening of a CDW gap in FeGe.
The kagome lattice has been a fruitful playground to study novel quantum phenomena for over 70 years, uniting electronic correlations, topologically nontrivial states, and geometric frustration [1, 2]. Being rooted in the network of corner-sharing triangles, destructive quantum interference results in localized states, i.e., a flat band over the entire Brillouin zone [3, 4]. Together with the van Hove singularities at the point, this can lead to exotic electronic orders, including flat-band ferromagnetism, Pomeranchuck instabilities, charge and spin bond order, unconventional superconductivity, nematicity, and Wigner crystallization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
Several of these intriguing phases have been identified in kagome metals, most prominently the exotic charge-density-wave (CDW) in the superconducting VSb ( = K, Rb, Cs) family [15, 16, 17, 7, 18, 12]. Recently, charge order in the context of a magnetic kagome metal has been suggested [19, 20, 21]. Hexagonal FeGe crystallizes in the space group featuring magnetic Fe-kagome planes stabilized by Ge1 atoms. Within one unit cell, these kagome layers are separated by a non-magnetic honeycomb lattice of Ge2 atoms as visualized in Fig. 1 (a). At room temperature, an out-of-plane antiferromagnetic (AFM) arrangement of the Fe-kagome layers is found, with ferromagnetically ordered Fe-moments within the kagome planes. This AFM order persists down to approximately 60 K, below which a canted AFM arrangement with an additional reorientation at 30 K is observed [22, 23].
More recent studies reveal an anomaly at K, where a slight drop in the in-plane magnetic susceptibility occurs [19, 24, 25]. However, no clear counterpart to this drop is observed in the electric resistivity [24, 25]. Based on neutron diffraction and scanning tunneling microscopy (STM) measurements, a short-range ordered CDW phase with a minor in-plane lattice distortion on the order of of the kagome structure has been proposed [19, 21, 24]. Raman spectroscopy and neutron Larmor diffraction further suggested that deviations from the hexagonal symmetry occur above [26].
Although the low-temperature, putative CDW phase of FeGe was investigated extensively by surface-sensitive techniques such as STM and angle-resolved photoemission spectroscopy (ARPES), a systematic study of the temperature-dependent bulk electronic structure remains lacking. Fourier-transform infrared spectroscopy is one of the most effective approaches for investigating bulk CDW order [27, 28, 29, 30, 31, 32, 33]. Here, the gap opening at the Fermi level below is normally manifested in a spectral weight transfer from low to high energies [34]. However, our results on FeGe reveal no indication of a gap opening below the anomaly at and, thus, contradict the CDW scenario at low temperatures. Aided by density functional theory (DFT) calculations of the band structure and the optical conductivity, we show that the reconstruction of the bands due to the structural transition at leads to new interband optical transitions at low energies. We thus interpret the change in the electronic structure below as band splitting rather than gap opening.
High-quality single crystals were synthesized as explained in the Supplemental Material [36]. Prior to the optical experiments, we measured the in-plane magnetic susceptibility and dc resistivity of our samples. The results are plotted in Fig. 1 (b), confirming the stoichiometry and identifying the low-temperature anomaly at K. Fig. 1 (c) displays the temperature-dependent real part of the in-plane optical conductivity. At first glance, the spectra resemble the optical spectra of other magnetic kagome metals such as the ferromagnetic FeSn [37, 38], and the ferrimagnetic MnSn compounds [39].
Using the Drude-Lorentz approach, different contributions to the total optical spectra are modeled. In addition to the classical Lorentzian (interband) and Drude (intraband) contributions, a sharp phonon mode around 190 cm with a Fano-like shape, and a temperature-dependent absorption peak at low energies are observed. The latter feature is assigned to the intraband signature of localized electrons, in line with previous optical studies of kagome metals [40, 41, 42, 43, 37, 39]. The total complex optical conductivity [] then takes the form
After decomposing the optical data, as shown in Fig. S2 for selected temperatures, it becomes apparent that the low-energy interband transitions strongly vary with temperature. To further visualize these changes, we subtract the Fano resonance and the intraband contributions, i.e., Drude and localization peaks from the optical conductivities as presented in Fig. 2 (a) and (b) for . The room temperature data is reproduced by our DFT calculations given in panels (c) and (d) via raising the Fermi level by 110 meV. Upon cooling, the characteristic changes in the experimental spectra can be described by a systematic reduction of the Fermi level. This finding is consistent with the ARPES study in Ref. [19] reporting a significant shift of the Fermi level at temperatures above . Aided by band-resolved optical conductivity calculations given in the Supplemental Material [36], the variation of the low-energy spectral weight can be related to changes of the transition energy between multiple linear Fe 3-bands along , , and [see Fig. 2 (e)]. Note that, for all calculations, the energy axis needs to be rescaled by 1/2.1 to achieve a good match with the experiment. A comparable rescaling factor of 1/1.6 was observed in ARPES measurements [20]. Moreover, a similar discrepancy between theory and experiment is reported for other magnetic kagome metals, such as the MnSn family and CoSnS, signaling a significant influence of electronic correlations [39, 44]. This effect is also observable in the difference between the experimental and DFT plasma frequencies, which can be used to gauge the strength of the electronic correlations as done in Fig. S8. Our data evidence much stronger correlations in FeGe compared to the nonmagnetic kagome metals, VSb and ScVSn, and even compared to the magnetic CoSnS.
Upon cooling below , several modifications in the electronic structure are clearly detected. First, the Fano resonance shifts to lower energies and significantly broadens upon approaching , indicating strong electron-phonon coupling. This behavior is highly unusual for the temperature evolution of phonon modes, which are expected to sharpen and blueshift as the lattice hardens with decreasing temperature (see Supplemental Material for a more detailed discussion on the Fano resonance [36]). Second, there is a notable enhancement in the spectral weight around 250 cm and a gradual formation of a sharp absorption feature at 100 cm, as displayed in Fig. 2 (f). However, no spectral weight transfer from low to high energies, i.e., no optical signature of a CDW gap, is observed. The absence of a gap opening in the optical spectra is consistent with the results of the Raman studies [45].
To further investigate the effect of lattice distortion below on the electronic structure, we carry out DFT calculations of the band structure and optical conductivity for the low-temperature phase. Below , a superstructure as depicted in Fig. 1 (a), involving a tri-hexagonal (trh) distortion of the Fe-kagome planes was observed by X-ray diffraction (XRD) measurements [19]. Additionally, a partial dimerization of Ge1 atoms was proposed by XRD and inelastic X-ray scattering studies on annealed samples, as well as DFT simulations [46, 47, 24, 48, 25]. In both cases, however, the distortion of the kagome planes is on the order of Å and, thus, extremely small.
The minuscule Fe displacement associated with the tri-hexagonal distortion induces only a tiny splitting between the parallel bands crossing the Fermi level (see Fig. S7 for a more detailed view). Thus, plenty of spectral weight is found at low energies, resulting in the pronounced peak at 100 cm as displayed in Fig. 3 (b). The splitting between some parallel bands can be increased by imposing a larger displacement of the Fe atoms. As discussed in the Supplemental Material [36], this leads to more spectral weight at energies between 200 and 1000 cm, with the low-energy peak at 100 cm persisting. However, there is still no indication of a gap opening neither in the calculated optical conductivity nor in the density of states, as shown in Fig. S6. Hence, our study suggests a different nature of the low-temperature transition in FeGe. Moreover, experimental manifestations of this transition appear to be dependent on the growth method and require sensitive measurements as it was not detected in early studies [49, 22]. Doping studies show a rapid suppression of the low-temperature anomaly in the magnetic susceptibility already at 1.5 % Sb-doping of the Ge-sites, militating in favor of the significant involvement of the Ge atoms in the low-temperature phase transition [50].
The Ge dimerization has a more drastic effect on the band structure compared to the deformation of the kagome planes. As seen in Fig. 2 (e) and Fig. 3 (c), precisely one in-plane Ge1 band crosses the Fermi level in the normal state. This band becomes split in the dimerized state with a gap of around 0.7 eV [panel (d)]. While contributions from this splitting to the optical data are masked by several other interband optical transitions due to the multi-band nature of FeGe, a reorganization of the Fe bands near leads to additional spectral weight at approximately 250 cm, consistent with the experimental result. Along with the improved agreement between theory and the experiment above 1000 cm, these findings support the Ge dimerization scenario. Furthermore, the involvement of Ge1 atoms may also explain why the low-temperature transition has rather strong signatures in the optical spectra, i.e., Fano anomalies and different interband transitions, even though almost no changes happen in the Fe-kagome planes.
The low-temperature anomaly in the magnetic susceptibility and the specific heat at can be enhanced by a post-growth annealing process. However, the signature of the corresponding phase transition in the electric resistivity remains astonishingly weak [24, 25, 45]. While additional Bragg peaks below imply a structural modulation [19], this does not necessarily lead to a CDW. Kagome systems are prone to multiple structural instabilities, such as Wigner crystallization [14, 9]. Moreover, breathing distortion of the kagome lattice is a well-known effect in quantum magnetism, but also in the kagome metals where FeSn adopts the breathing structure without any CDW [51, 52, 53, 54].
Overall, there is a sharp contrast between FeGe and other kagome metals exhibiting a CDW at low temperatures, including the renowned VSb compounds. Their kagome planes exhibit a very similar structural distortion at low temperatures, forming a 2 2 in-plane superstructure. However, being on the order of Å, the atomic displacement is much more pronounced in the VSb compounds [55]. Interestingly, in their normal state, all these kagome metals feature multiple -bands and exactly one -band crossing the Fermi level. This -band is related to the in-plane Ge1 atoms in FeGe but to out-of-plane Sb2 atoms in VSb [56]. It essentially remains intact in the CDW phase of VSb [55] but splits in FeGe due to the partial dimerization of the in-plane Ge1 atoms. Lastly, our data do not reveal the presence of a gap in the low-temperature phase of FeGe, while clear signatures of a gap opening in ScVSn and the VSb compounds were detected by infrared spectroscopy [40, 41, 42, 57, 58, 59, 60].
All these findings imply that the low-temperature phase of FeGe is substantially different from the CDW phase observed in other kagome metals. On the one hand, our study supports the partial dimerization of the in-plane Ge atoms, being decisive for the significant changes in the optical data below the anomaly at . On the other hand, the absence of a gap in both experiment and theory is clearly different from the conventional CDW scenario that has been discussed for FeGe previously. Given the complex interplay between magnetism, structural instabilities, strong electronic correlations, as well as the simple tunability of the low-temperature anomaly by doping and annealing [50, 45, 24, 25], antiferromagnetic FeGe might open new routes to study so far unexplored regions in the rich phase diagram of kagome metals.
Authors acknowledge the fruitful discussions with S. Fratini and the technical support by G. Untereiner. M.W. is supported by IQST Stuttgart/Ulm via a project funded by the Carl Zeiss Stiftung. The work has been supported by the Deutsche Forschungsgemeinschaft (DFG) via DR228/51-3, DR228/68-1, and UY63/2-1.